OpenStudy (anonymous):

The derivative of 3/t^3/5 - 6/t^1/2

OpenStudy (mathteacher1729):

Do you mean \[\frac{dy}{dx}(\frac{3}{t^{3/5}}-\frac{6}{t^{1/2}})\] ?

OpenStudy (anonymous):

Yes

OpenStudy (mathteacher1729):

Ok, you must use two rules here: \[1/x^n = x^{-n}\] and \[\frac{dy}{dx}x^k = kx^{k-1}\] So re-write this as \[\frac{dy}{dx}(3t^{-3/5}-6t^{-1/2})\]

OpenStudy (anonymous):

Thanks!

OpenStudy (mathteacher1729):

You do know that what I wrote is not the final answer, right? You still have to differentiate according to those rules. :)